OpenStudy (anonymous):

Find the ninth term of this sequence: 81/8, 27/4, 9/2, 3,...

sam (.sam.):

what do you think about the sequence, ap? gp?

OpenStudy (anonymous):

whats ap, gp?

sam (.sam.):

ap, arithmetic progression gp, geometric progression

OpenStudy (anonymous):

Ooo im not sure about that stuff

sam (.sam.):

try to find the difference between the numbers for ap, second one minus the first one, for gp, second one divide the first one

sam (.sam.):



OpenStudy (anonymous):

oh okaay

sam (.sam.):

for gp use \[t_n=ar^{n-1}\] since its ninth term, a=81/8, r=3/2 \[t_9=(81/8)(3/2)^{9-1}\] solve

OpenStudy (anonymous):

is that a g next to the t?

sam (.sam.):

its 9

OpenStudy (anonymous):

oh ok sorry its a little blurry to me on my computer.

OpenStudy (anonymous):

i got 259.4..